# Generative Inverse Design of Metamaterials with Functional Responses by Interpretable Learning

### v3

@@ 1 Introduction
Metamaterials with functional responses are engineered materials that exhibit varying properties or behaviors under different conditions. One example is metamaterials whose electromagnetic, acoustic, or elastic wave propagation behaviors change with wavelengths or frequencies [1]. Another example is metamaterials that exhibit changing properties or functionalities due to deformation in response to external stimuli like temperature [2] and magnetic fields [3]. Tailoring functional responses of these metamaterials is of interest to applications such as sound and vibration control, analog computing, medical imaging, sensing, communication, and soft robotics. In many use cases, rather than precisely controlling the complete functional responses, we only care about qualitative behaviors under certain conditions. For example, acoustic metamaterials were usually designed to have bandgaps at specified frequencies to achieve functionalities like wave-guiding [4, 5], focusing [6, 7], and vibration mitigation [8, 9, 10]. However, it is unnecessary and computationally expensive to design for the whole dispersion relation [11, 12, 13, 14, 15]. Similarly, we may design optical metamaterials to qualitatively manipulate optical properties (e.g., high or low absorption/reflection/transmission) under certain wavelengths, without requiring the entire spectral response to match an exact target [16, 17].
@@
Prior research attempted to realize iteration-free, single-shot inverse design using machine learning. There are three mainstream methods. The most straightforward approach is to learn a direct inverse mapping from the response to design variables. Neural networks are the most commonly used machine learning model for this purpose, due to their high flexibility in approximating arbitrary nonlinear input-output relationships [25, 13]. Despite the simplicity of the direct inverse mapping, its underlying assumption of the response-design mapping being one-to-one does not hold in many cases due to the non-uniqueness of solutions, as mentioned earlier. Such non-uniqueness will cause conflicting training instances where the same input (response) is associated with distinct outputs (designs), which will destabilize the convergence during neural network training [22, 26]. To avoid this issue, past work proposed the Tandem Neural Network (T-NN) that cascades an inverse-design network with a pretrained forward-modeling network to avoid using designs as labels when training the inverse-design network and hence solved the training convergence issue [22, 27, 28, 29, 30, 31, 32]. Nonetheless, the original T-NNs still learn a one-to-one response-design mapping and cannot account for the non-uniqueness of design solutions. To fundamentally solve this problem, one needs to learn a one-to-many mapping. For example, Bastek et al. [30] integrated stochastic sampling into the inverse-design network to allow the generation of multiple feasible solutions. Ha et al. [33] used multiple T-NNs to predict multiple design solutions given a target response. Wang et al. [34] adapted the T-NN to mimic the structure of a variational autoencoder that is capable of learning the conditional distributions of designs given target responses. A large body of recent works achieved the goal of learning one-to-many mapping by using conditional generative models, typically conditional generative adversarial networks (cGANs) [35, 36, 37, 38, 24], conditional variational autoencoders (cVAEs) [23, 39, 40], and conditional diffusion models [41, 42, 43]. These models can generate multiple designs given a target response by learning the distribution of designs conditioned on the response. We include more detailed explanations of these three mainstream methods in SI Appendix, Sec. A and Fig. S1.
@@ 2 Methods
Figure 2 shows how, by probing the trained random forest, one can estimate a likelihood distribution for target satisfaction of solutions over the entire design space and sample (generate) new designs based on this likelihood distribution. Since we target qualitative (binary) behaviors at specified $s$ (e.g., a bandgap in 3-4 MHz frequency or high absorption at a wavelength of 400-500 nm), we first identify the leaf nodes (on each decision tree in the random forest) that are relevant to the $s$ in the target (Fig. 2, Step 1). We do this by tracing down each tree, checking only the nodes that use $s$ as the splitting feature, and pruning the branches that are irrelevant to the $s$ in the target. For example, as shown in Fig. 2, there are two tree nodes using $s$ as the splitting feature, with splitting criteria at $s\leq 5$ and $s\leq 7$. Given the target frequency range of $3\leq s\leq 4$, we can remove the right branches of both nodes as these branches are only relevant to $s>5$ and $s>7$, respectively, which conflicts with the target range of $3\leq s\leq 4$. After pruning these branches, we end up with a set of leaves relevant to the target (highlighted in Fig. 2, Step 1). When we have a combined target (e.g., bandgaps in both 3-4 MHz and 6-7 MHz, as shown in Fig. 2), we need to get the intersection of all the sets of relevant leaves and use that as the final set of relevant leaves (highlighted in Fig. 2, Step 2). Note that a combined target includes cases where there are multiple nonadjacent target ranges (e.g., 3-4 MHz and 6-7 MHz) or when a target range is split by a tree node (e.g., a target range of 4-6 MHz can be split by the node “$s\leq 5$”, thus we need to consider it as the combination of two target ranges — 4-5 MHz and 5-6 MHz). A more detailed discussion of this step is in SI Appendix, Sec. B.
@@
Since a single decision tree usually lacks accuracy, robustness, and a way to quantify estimation uncertainty, we still want to take advantage of the random forest as an ensemble model for inverse design. We use Steps 1-3 to derive the likelihood distribution for each of the $M$ trees in the random forest, and simply use the average of these $M$ likelihood distributions as the final likelihood for target satisfaction, $\mathcal{L}(\mathbf{x}|\mathcal{T})=\sum_{m}^{M}\mathcal{L}_{m}(\mathbf{x}|\mathcal{T})/M$, which is a more complex and smooth function (Fig. 2, Step 4). If more trees believe a design $\mathbf{x}$ has a higher likelihood of satisfying the target, then the design will have a higher likelihood $\mathcal{L}(\mathbf{x}|\mathcal{T})$. Finally, to generate new designs, we can sample from $\mathcal{L}(\mathbf{x}|\mathcal{T})$ using Markov chain Monte Carlo (MCMC) methods such as Metropolis-Hastings [52] (Fig. 2, Step 5).
@@ 3 Results
For all the test problems, we used the same random forest hyperparameter settings and did not perform hyperparameter tuning. Specifically, each random forest contains 1,000 trees with a minimum of 2 samples required to split an internal node and a minimum of 1 sample required to be at a leaf node. Gini impurity [53] was used as the splitting criterion at tree nodes. The train-test split ratio was 4:1. Since the positive/negative training data can be highly imbalanced (e.g., the frequency ranges with bandgaps are much narrower than those without), we used the Synthetic Minority Over-sampling TEchnique (SMOTE) [54] to over-sample the positive class. For all the case studies, the random forest training took less than 10 seconds on an Intel Core i5-9300H CPU 2.4GHz and 8GB memory.
After training, we generate new designs by sampling from the resulting likelihood distribution using Metropolis-Hastings. In practice, Metropolis-Hastings can generate identical samples, which provides no benefits for design purposes. Thus in this work, we reject the designs identical to previous ones during sampling.
@@ 3.1 Applying RIGID to Design Acoustic Metamaterials with Target Bandgaps
Figure 3: Acoustic metamaterial design problem configuration and results. (A) Design variables of center and corner mass radii ($r_{\text{center}}$ and $r_{\text{corner}}$) and strut radius ($r_{\text{strut}}$). (B) High symmetry points of the cubic irreducible Brillouin zone. (C) A sample dispersion relation and bandgap (marked by the highlighted zone). The design objective is to generate new acoustic metamaterial designs with target bandgaps. (D) KDE of the estimated likelihood for generated designs. (E) Validation metrics under varying sampling thresholds. (F) Geometries and corresponding dispersion relations of five generated designs with the highest likelihood of satisfying a specified target bandgap (marked as highlighted frequency regions). All the radii ($r_{\text{strut}}$, $r_{\text{center}}$, and $r_{\text{corner}}$) have a unit of $\mu$m. Here only the fourth design fails to meet a small portion (at around 6 MHz) of the target bandgap, whereas the others meet it. Generated designs for other targets are shown in SI Appendix, Figs. S2-S6.
@@
To test the inverse design capability of RIGID, we randomly created 10 targets, each containing 1-2 frequency ranges in which bandgap(s) should exist. We generated 30 designs for each target by sampling from the resulting likelihood distribution over the design space [^1] Note that it is possible for the likelihood to be zero everywhere in the design space when the model believes the target is unachievable. We ignore these cases as it is meaningless and impossible to sample designs from such likelihood distribution.. Bandgaps were identified from dispersion relations computed using Bloch-wave analysis. Figure 3D shows the kernel density estimation (KDE) for the likelihood of the 300 generated designs, conditioned on their target satisfaction. We use $\mathcal{D}$ and $\mathcal{D}_{\text{feas}}$ to represent the complete set of generated designs and the set of generated designs that actually satisfy the target, respectively. Then $\mathcal{D}\backslash\mathcal{D}_{\text{feas}}$ denotes the set of generated designs that cannot fulfill the target in reality.
In an ideal scenario, all solutions in $\mathcal{D}$ would satisfy the target, which means $\mathcal{D}=\mathcal{D}_{\text{feas}}$, and their density profiles should coincide. However, this ideal scenario is impossible in reality due to both limited model accuracy and uncertainty. As RIGID generates designs by sampling the entire design space based on the likelihood values, it is expected for non-satisfactory designs to be generated as long as the likelihood values for these designs are non-zeros. Therefore, the fact that a design is generated does not mean that the model is certain about the design satisfying the target. We have to look at the estimated likelihood to know how likely the target will be satisfied and make decisions from there. RIGID thus incorporated the model’s confidence into the inverse design process.
For a reasonable model, most designs in $\mathcal{D}\backslash\mathcal{D}_{\text{feas}}$ should have low estimated likelihood values. Consequently, the density of $\mathcal{D}_{\text{feas}}$’s likelihood is a result of shifting some of $\mathcal{D}$’s density from left (low likelihood) to right (high likelihood). This expectation aligns with the observation in Fig. 3D.
@@
Figure 3F visualizes the geometries and dispersion relations of designs generated based on a randomly created bandgap target. Only the top five designs with the highest likelihood values are shown. In this example, our method generates geometrically different designs that have a high probability of achieving target bandgaps, each yielding a slightly different dispersion relation. This is promising in design applications requiring other material properties, such as dynamic wave velocity or quasi-static stiffness, in which the user can select from a menu of designs with the same target bandgap but other varying properties. Generated designs based on the other nine bandgap targets can be found in SI Appendix, Figs. S2-S6.
@@ 3.2 Applying RIGID to Design Optical Metasurfaces with Target High-Absorbance Wavelengths
Figure 4: Optical metasurface design problem configuration and results. (A-B) Design variables (materials, layer thicknesses, and cross-section geometry types). (C) A sample absorbance spectrum and the wavelength intervals (highlighted wavelength regions) corresponding to absorbance above the threshold $t$. The design objective is to generate new optical metasurface designs that exhibit higher absorbance than a threshold $t$ at the user-defined wavelength interval(s). (D) KDE of the estimated likelihood for generated designs. (E) Validation metrics under varying sampling thresholds. (F) Designs (geometries and material selections) and corresponding absorbance spectra of five generated metasurfaces with the highest likelihood of satisfying specified target high-absorbance regions (marked as highlighted wavelength regions). All the layer thicknesses ($h_{l},l=1,2,3$) have a unit of nm. Here all five designs satisfy the target. Generated designs for other targets are shown in SI Appendix, Figs. S7-S9.
@@
Figure 4F shows generated optical metasurface designs with the top five likelihood estimations for a randomly created target. While the materials, cross-section geometries, and layer thicknesses of generated designs can be different, all the designs satisfy the target (Fig. 4F, right panel). To further enhance the diversity of final solutions, we can use sampling strategies such as the one proposed in Ref. [76, 77] to identify a subset (of generated solutions) that simultaneously exhibits high likelihood and high diversity. Generated designs based on the other nine targets can be found in SI Appendix, Figs. S7-S9.

### v4

@@ 1 Introduction
Metamaterials are engineered materials, typically composed of periodically arranged building blocks (i.e., unit cells), that exhibit properties and functionalities beyond their constituent materials [1, 2]. We can achieve certain effective material properties of metamaterials by designing their structure (i.e., geometries of unit cells) rather than chemical composition. Optical [3], acoustic [4], thermal [5], and mechanical metamaterials [6, 2] are metamaterials with unique properties tailored to manipulate specific types of waves or energy. For example, optical metamaterials or metasurfaces (i.e., the two-dimensional form of metamaterials) are designed to control and manipulate electromagnetic waves, which can lead to unusual properties such as a negative refractive index [7]. Acoustic metamaterials are designed to control how sound (acoustic) waves propagate which enables functionalities such as noise reduction [8].
Functional responses refer to varying properties or behaviors under different conditions. One example is metamaterials whose electromagnetic, acoustic, or elastic wave propagation behaviors change with wavelengths or frequencies [9]. Another example is metamaterials that exhibit changing properties or functionalities due to deformation in response to external stimuli like temperature [10] and magnetic fields [11]. Tailoring functional responses of these metamaterials is of interest to applications such as sound and vibration control, analog computing, medical imaging, sensing, communication, and soft robotics. In many use cases, rather than precisely controlling the complete functional responses, we only care about qualitative behaviors under certain conditions. For example, acoustic metamaterials were usually designed to have bandgaps at specified frequencies to achieve functionalities like wave-guiding [12, 13], focusing [14, 15], and vibration mitigation [16, 17, 18]. However, it is unnecessary and computationally expensive to design for the whole dispersion relation [19, 20, 21, 22, 23]. Similarly, we may design optical metamaterials to qualitatively manipulate optical properties (e.g., high or low absorption/reflection/transmission) under certain wavelengths, without requiring the entire spectral response to match an exact target [24, 25].
@@
Prior research attempted to realize iteration-free, single-shot inverse design using machine learning. There are three mainstream methods. The most straightforward approach is to learn a direct inverse mapping from the response to design variables. Neural networks are the most commonly used machine learning model for this purpose, due to their high flexibility in approximating arbitrary nonlinear input-output relationships [33, 21]. Despite the simplicity of the direct inverse mapping, its underlying assumption of the response-design mapping being one-to-one does not hold in many cases due to the non-uniqueness of solutions, as mentioned earlier. Such non-uniqueness will cause conflicting training instances where the same input (response) is associated with distinct outputs (designs), which will destabilize the convergence during neural network training [30, 34]. To avoid this issue, past work proposed the Tandem Neural Network (T-NN) that cascades an inverse-design network with a pretrained forward-modeling network to avoid using designs as labels when training the inverse-design network and hence solved the training convergence issue [30, 35, 36, 37, 38, 39, 40]. Nonetheless, the original T-NNs still learn a one-to-one response-design mapping and cannot account for the non-uniqueness of design solutions. To fundamentally solve this problem, one needs to learn a one-to-many mapping. For example, Bastek et al. [38] integrated stochastic sampling into the inverse-design network to allow the generation of multiple feasible solutions. Ha et al. [41] used multiple T-NNs to predict multiple design solutions given a target response. Wang et al. [42] adapted the T-NN to mimic the structure of a variational autoencoder that is capable of learning the conditional distributions of designs given target responses. A large body of recent works achieved the goal of learning one-to-many mapping by using conditional generative models, typically conditional generative adversarial networks (cGANs) [43, 44, 45, 46, 32], conditional variational autoencoders (cVAEs) [31, 47, 48], and conditional diffusion models [49, 50, 51]. These models can generate multiple designs given a target response by learning the distribution of designs conditioned on the response. We include more detailed explanations of these three mainstream methods in SI Appendix, Sec. S1 and Fig. S1.
@@ 2 Methods
Figure 2 shows how, by probing the trained random forest, one can estimate a likelihood distribution for target satisfaction of solutions over the entire design space and sample (generate) new designs based on this likelihood distribution. Since we target qualitative (binary) behaviors at specified $s$ (e.g., a bandgap in 3-4 MHz frequency or high absorption at a wavelength of 400-500 nm), we first identify the leaf nodes (on each decision tree in the random forest) that are relevant to the $s$ in the target (Fig. 2, Step 1). We do this by tracing down each tree, checking only the nodes that use $s$ as the splitting feature, and pruning the branches that are irrelevant to the $s$ in the target. For example, as shown in Fig. 2, there are two tree nodes using $s$ as the splitting feature, with splitting criteria at $s\leq 5$ and $s\leq 7$. Given the target frequency range of $3\leq s\leq 4$, we can remove the right branches of both nodes as these branches are only relevant to $s>5$ and $s>7$, respectively, which conflicts with the target range of $3\leq s\leq 4$. After pruning these branches, we end up with a set of leaves relevant to the target (highlighted in Fig. 2, Step 1). When we have a combined target (e.g., bandgaps in both 3-4 MHz and 6-7 MHz, as shown in Fig. 2), we need to get the intersection of all the sets of relevant leaves and use that as the final set of relevant leaves (highlighted in Fig. 2, Step 2). Note that a combined target includes cases where there are multiple nonadjacent target ranges (e.g., 3-4 MHz and 6-7 MHz) or when a target range is split by a tree node (e.g., a target range of 4-6 MHz can be split by the node “$s\leq 5$”, thus we need to consider it as the combination of two target ranges — 4-5 MHz and 5-6 MHz). A more detailed discussion of this step is in SI Appendix, Sec. S2.
@@
Since a single decision tree usually lacks accuracy, robustness, and a way to quantify estimation uncertainty, we still want to take advantage of the random forest as an ensemble model for inverse design. We use Steps 1-3 to derive the likelihood distribution for each of the $M$ trees in the random forest, and simply use the average of these $M$ likelihood distributions as the final likelihood for target satisfaction, $\mathcal{L}(\mathbf{x}|\mathcal{T})=\sum_{m=1}^{M}\mathcal{L}_{m}(\mathbf{x}|\mathcal{T})/M$, which is a more complex and smooth function (Fig. 2, Step 4). If more trees believe a design $\mathbf{x}$ has a higher likelihood of satisfying the target, then the design will have a higher likelihood $\mathcal{L}(\mathbf{x}|\mathcal{T})$. Finally, to generate new designs, we can sample from $\mathcal{L}(\mathbf{x}|\mathcal{T})$ using Markov chain Monte Carlo (MCMC) methods such as Metropolis-Hastings [60] (Fig. 2, Step 5).
We can also derive the posterior of a design $\mathbf{x}$ given the target if we know the prior $p(\mathbf{x})$ based on Bayes’ theorem:
$p(\mathbf{x}|\mathcal{T})\propto\mathcal{L}(\mathbf{x}|\mathcal{T})p(\mathbf{x})$
In this work, we assume a uniform prior, which is equivalent to just considering the likelihood $\mathcal{L}(\mathbf{x}|\mathcal{T})$.
@@ 3 Results
For all the test problems, we used the same random forest hyperparameter settings and did not perform hyperparameter tuning. Specifically, each random forest contains 1,000 trees with a minimum of 2 samples required to split an internal node and a minimum of 1 sample required to be at a leaf node. Gini impurity [61] was used as the splitting criterion at tree nodes. The train-test split ratio was 4:1. Since the positive/negative training data can be highly imbalanced (e.g., the frequency ranges with bandgaps are much narrower than those without), we used the Synthetic Minority Over-sampling TEchnique (SMOTE) [62] to over-sample the positive class. For all the case studies, the random forest training took less than 10 seconds on an Intel Core i5-9300H CPU 2.4GHz and 8GB memory. After training, we generate new designs by sampling from the resulting likelihood distribution using Metropolis-Hastings. In practice, Metropolis-Hastings can generate identical samples, which provides no benefits for design purposes. Thus in this work, we reject the designs identical to previous ones during sampling.
We compare RIGID to a commonly used low-cost generative design strategy—surrogate modeling combined with the genetic algorithm (GA) [63, 64, 65, 66, 67]. We choose this baseline according to the following criteria to ensure that its efficiency and application scenario are comparable to those of RIGID: 1) the baseline method should generate design solutions quickly (i.e., no additional training or design evaluation is required during the use phase), 2) the baseline method can generate multiple satisfying solutions to the inverse design problem, and 3) the baseline method can handle small data. These criteria exclude sequential design strategies like Bayesian optimization (as they violate the first criterion) and any existing single-shot inverse design methods mentioned in Sec. 1 (as direct inverse mapping and T-NN violate the second criterion and conditional generative model-based methods usually violate the last criterion).
The genetic algorithm is an optimization method that iteratively evolves a population of candidate solutions, based on biologically inspired operators such as mutation, crossover, and selection, toward better solutions [68]. As we usually need to evaluate a large population of candidate solutions at each iteration (referred to as a generation), a surrogate model (usually a machine learning model) is often constructed to accelerate design response prediction, thereby speeding up the optimization process. For the surrogate model in the baseline method, we use the same trained random forest as in RIGID for consistency, eliminating differences due to forward prediction accuracy and focusing on comparing the inverse design methods. We set the objective of GA to maximize the aggregated probability of target satisfaction within the target domain $\Omega$, i.e., $\max_{\mathbf{x}}\sum_{s_{i}\in\Omega}f^{\prime}(\mathbf{x},s_{i})$, where $\{s_{i}\}_{1}^{n}$ are discretized auxiliary variable values (e.g., discrete values of frequencies or wavelengths) and $f^{\prime}$ represents the trained random forest. The initial population for GA is generated by randomly sampling solutions within the design variable bounds. At each generation, we select individuals to reproduce for a new generation by randomly picking three individuals from the population, selecting the best one, and repeating this process 300 times. We set the probability for crossover and mutation to be 0.5 and 0.2, respectively. Once an individual is mutating, there is a 5% chance for each design variable in this individual to mutate. For a continuous design variable, we define the mutation operation as adding Gaussian noise with a mean of 0 and a standard deviation equal to 1% of the variable’s range. For a categorical design variable, mutation is defined as randomly reassigning the category of the variable. We ran GA for 10 generations, each with a population of 300. Finally, we select a set of solutions with the highest objective values among all the individuals created during GA, and compare them to designs generated by RIGID. From this point, we will call designs generated by RIGID and GA as RIGID designs and GA designs, respectively.
@@ 3.1 Applying RIGID to Design Acoustic Metamaterials with Target Bandgaps
Figure 3: Acoustic metamaterial design problem configuration and results. (A) Design variables of center and corner mass radii ($r_{\text{center}}$ and $r_{\text{corner}}$) and strut radius ($r_{\text{strut}}$). (B) High symmetry points of the cubic irreducible Brillouin zone. (C) A sample dispersion relation and bandgap (marked by the highlighted zone). The design objective is to generate new acoustic metamaterial designs with target bandgaps. (D) KDE of the estimated likelihood for generated designs. (E) Satisfaction rates, average scores, and selection rates for RIGID designs under varying sampling thresholds (solid lines), in comparison to the satisfaction rate of GA designs (horizontal dashed line). The horizontal dotted line indicates 100% satisfaction (F) Geometries and corresponding dispersion relations of five RIGID designs with the highest likelihood of satisfying a specified target bandgap (marked as highlighted frequency regions). All the radii ($r_{\text{strut}}$, $r_{\text{center}}$, and $r_{\text{corner}}$) have a unit of $\mu$m. Here only the fourth design fails to meet a small portion (at around 6 MHz) of the target bandgap, whereas the others meet it. Generated designs for other targets are shown in SI Appendix, Figs. S2-S6.
@@
To test the inverse design capability of RIGID, we randomly created 10 targets, each containing 1-2 frequency ranges in which bandgap(s) should exist. We generated 30 designs for each target by sampling from the resulting likelihood distribution over the design space [^1] Note that it is possible for the likelihood to be zero everywhere in the design space when the model believes the target is unachievable. We ignore these cases as it is meaningless and impossible to sample designs from such likelihood distribution.. Bandgaps were identified from dispersion relations computed using Bloch-wave analysis.
Figure 3D shows the kernel density estimation (KDE) for the likelihood of the 300 RIGID designs, conditioned on their target satisfaction. We use $\mathcal{D}$ and $\mathcal{D}_{\text{feas}}$ to represent the complete set of generated designs and the set of generated designs that actually satisfy the target, respectively. Then $\mathcal{D}\backslash\mathcal{D}_{\text{feas}}$ denotes the set of generated designs that cannot fulfill the target in reality. In an ideal scenario, all solutions in $\mathcal{D}$ would satisfy the target, which means $\mathcal{D}=\mathcal{D}_{\text{feas}}$, and their density profiles should coincide. However, this ideal scenario is impossible in reality due to both limited model accuracy and uncertainty. As RIGID generates designs by sampling the entire design space based on the likelihood values, it is expected for non-satisfactory designs to be generated as long as the likelihood values for these designs are non-zeros. Therefore, the fact that a design is generated does not mean that the model is certain about the design satisfying the target. We have to look at the estimated likelihood to know how likely the target will be satisfied and make decisions from there. RIGID thus incorporated the model’s confidence into the inverse design process. For a reasonable model, most designs in $\mathcal{D}\backslash\mathcal{D}_{\text{feas}}$ should have low estimated likelihood values. Consequently, the density of $\mathcal{D}_{\text{feas}}$’s likelihood is a result of shifting some of $\mathcal{D}$’s density from left (low likelihood) to right (high likelihood). This expectation aligns with the observation in Fig. 3D.
@@
In the comparative study, we generated the same number of designs for the same targets by applying GA. We computed the satisfaction rate of GA designs and found that the value is close to the satisfaction rate of RIGID designs when setting $\tau$ to 0.2. This means that if we set $\tau$ to be above 0.2, RIGID will most likely achieve a higher satisfaction rate than GA. The distributions of generated satisfactory solutions shown in Fig. 4 indicate that GA designs are highly localized while RIGID can capture the high diversity of the inverse design solutions, and sometimes discover solutions very different from the original feasible solutions from data. Besides, Fig. 4 also shows an example where GA fails to find any satisfactory solution while RIGID can still generate a very diverse set of satisfactory solutions.
Figure 4: Distributions of satisfactory solutions for two bandgap targets. The left panel shows that GA designs are highly localized while RIGID can lead to diverse solutions. The right panel indicates that none of the GA designs satisfy the target, while satisfactory RIGID designs are diverse and can be very different from feasible designs from data. Solutions from data include feasible designs in both training and test data.
Figure 3F visualizes the geometries and dispersion relations of RIGID designs generated based on a randomly created bandgap target. Only the top five designs with the highest likelihood values are shown. In this example, our method generates geometrically different designs that have a high probability of achieving target bandgaps, each yielding a slightly different dispersion relation. This is promising in design applications requiring other material properties, such as dynamic wave velocity or quasi-static stiffness, in which the user can select from a menu of designs with the same target bandgap but other varying properties. Generated designs based on the other nine bandgap targets can be found in SI Appendix, Figs. S2-S6.
@@ 3.2 Applying RIGID to Design Optical Metasurfaces with Target High-Absorbance Wavelengths
Figure 5: Optical metasurface design problem configuration and results. (A-B) Design variables (materials, layer thicknesses, and cross-section geometry types). (C) A sample absorbance spectrum and the wavelength intervals (highlighted wavelength regions) corresponding to absorbance above the threshold $t$. The design objective is to generate new optical metasurface designs that exhibit higher absorbance than a threshold $t$ at the user-defined wavelength interval(s). (D) KDE of the estimated likelihood for generated designs. (E) Satisfaction rates, average scores, and selection rates for RIGID designs under varying sampling thresholds (solid lines), in comparison to the satisfaction rate of GA designs (horizontal dashed line). (F) Designs (geometries and material selections) and corresponding absorbance spectra of five metasurfaces generated by RIGID. These five solutions are generated designs with the highest likelihood of satisfying specified target high-absorbance regions (marked as highlighted wavelength regions). All the layer thicknesses ($h_{l},l=1,2,3$) have a unit of nm. Here all five designs satisfy the target. Generated designs for other targets are shown in SI Appendix, Figs. S7-S9. (G) Distributions of design variables for satisfactory solutions generated by RIGID and GA (for the same target defined in Panel F). GA designs are highly localized and lack diversity compared to RIGID designs.
@@
We also generated the same amount of designs under the same targets by using GA. We found that GA designs can generate 100% satisfaction rate (Fig. 5E). However, similar to the acoustic metamaterial design example, the generated designs are localized compared to RIGID designs which cover much wider feasible region(s) in the design space. This can be reflected by the design variable distributions of generated feasible solutions shown in Fig. 5G, which corresponds to generated designs for a randomly created target (the target is shown as the highlighted wavelength regions in Fig. 5F). For example, all the GA designs have the same cross-section geometric type while RIGID generated satisfactory designs with different geometric types. It is easy to see that, without considering solution diversity, achieving a high target satisfaction rate is trivial—as long as we discover one satisfactory solution, we can add small perturbations to that solution to generate an infinite number of alternative solutions that have similar responses and thus are also likely to satisfy the target. Note that compared to the acoustic problem, this optical problem is less challenging as there is an average of 28.5% data already satisfying the ten targets (in contrast to an average of 9.3% satisfactory designs from data in the acoustic problem). Therefore, finding a few satisfactory solutions for the optical problem is relatively easy; the challenge lies in discovering all feasible solutions, where RIGID excels by identifying a wide range of them. We can also see that when the sampling threshold is 0.8, RIGID achieves a satisfaction rate of 100%, demonstrating its ability to trade the diversity of generated designs for higher feasibility.
Figure 5F shows generated RIGID designs with the top five likelihood estimations for a randomly created target. While the materials, cross-section geometries, and layer thicknesses of generated designs can be different, all the designs satisfy the target. Generated designs based on the other nine targets can be found in SI Appendix, Figs. S7-S9.